### v1

@@ 3.1. Shelling Order
    If $\alpha_{j}\geq n-4$, then $j_{1},j_{2}>\alpha_{j}$ implies $W_{n}[F_{j}^{c}]$ is connected. So $\alpha_{j}\leq n-5$, and hence $\alpha_{i}\geq 2m-\alpha_{j}\geq 5$. We have $j_{2}>j_{1}>\alpha_{j}>m\geq\alpha_{i}\geq 5$. Thus $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. Now, if $j_{1}\nsim\alpha_{i}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{i}$ and $j_{2}\nsim\alpha_{j}$ imply that $j_{2}\nsim\alpha_{i}$).
    We now assume that $j_{1}\sim\alpha_{i}$. Suppose $j_{2}\sim\alpha_{i}$. Then $j_{1}<\alpha_{i}\leq m<\alpha_{j}<j_{2}$ implies that either $\alpha_{i}=m$, $\alpha_{j}=m+1$, $j_{2}=m+2$ and $j_{1}\in\{m-2,m-1\}$, or $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. In the former case, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{1})$, a contradiction. So, we have $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. Then $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$ implies that $\alpha_{j}=n-1=j_{2}$, which is again a contradiction. Therefore, $j_{2}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(W_{n}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(W_{n}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. Since $W_{n}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{2}\nsim\alpha_{j}$. Therefore, $j_{2}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we get $F^{\prime}\prec F_{j}$ by Proposition 3.5. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
We show that $F_{q}\in M(\Delta_{3}(W_{n}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
@@
    Since $m<\alpha_{i}<\alpha_{j}$, we have $\alpha_{j}\geq m+2$. Then $\alpha_{j}<_{\mathcal{O}}j_{2}<\alpha_{j}$ implies that $j_{2}\leq m-3$. It follows that $j_{1}<j_{2}\leq m-3<m<\alpha_{i}<\alpha_{j}$. Thus $j_{2}\nsim\alpha_{i}$. Further, since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    We have $j_{2}>j_{1}>\alpha_{j}>\alpha_{i}>m>1$. So $j_{2}\nsim\alpha_{i}$. Since $W_{n}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    Since $m<\alpha_{i}<\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}j_{1}<\alpha_{j}$, we get $j_{1}\leq m-3$. Hence $j_{1}\leq m-3<m<\alpha_{i}<\alpha_{j}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Now, since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(W_{n}))$ (as $j_{2}\nsim\alpha_{j}$ implies that $j_{2}\nsim\alpha_{i}$).
@@
Suppose that $F_{j}\in\mathcal{D}$. Then there exists an $F_{r}\in M(\Delta_{3}(W_{n}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$, ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ and $F_{i}\prec F_{j}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, i.e., $i_{0}=j_{0}$, then we have $\alpha_{i}=\alpha_{j}$. In this case, if $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then $F_{r}=F_{i}$ satisfies ($\ast$3.1). So we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$ whenever $F_{i}\in\mathcal{T}_{j_{0}}$. Moreover, if $F_{i}\notin\mathcal{T}_{j_{0}}$, then $F_{i}\prec F_{j}$ and $F_{j}\in\mathcal{D}$ imply that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by Definition 3.4. Note that since $n\geq 9$, we have $45\leq m\leq n-$ 4. Now, $F_{j}\in\mathcal{D}$ implies that $F_{j}$ satisfies one of the conditions from $(\mathcal{D}_{1})$ to $(\mathcal{D}_{4})$.
@@
    Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. First we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$.
    First, suppose $i_{1}<j_{1}$. Since $i_{1}<j_{1}=\alpha_{j}-3<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$. Therefore, $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii) and (iv). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Let $i_{2}>j_{2}$. Since $\alpha_{j}\geq 6$, $j_{1}=\alpha_{j}-3\geq 3$. So we have $3\leq j_{1}=\alpha_{j}-3<\alpha_{j}<\alpha_{j}+1=j_{2}<i_{2}$. It follows that $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Hence $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-3$ and $i_{2}\geq\alpha_{j}+2$, $F^{\prime\prime}\notin\mathcal{D}$. Therefore $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii), and thus $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Now, suppose $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=\alpha_{j}+1$ implies that $i_{2}<\alpha_{j}$. Therefore, we have $\alpha_{j}-3=j_{1}<i_{1}<i_{2}<\alpha_{j}$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We know that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$. Thus, $\alpha_{i}\in\{m-1,m\}$. We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Therefore, since $W_{n}[F_{i}^{c}]$ is disconnected, we get $\alpha_{i}=m-1$, $i_{1}=j_{1}=m-2$ and $i_{2}=j_{2}=m+2$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
    Suppose $j_{1}\leq i_{1}<i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then we have $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. This implies that $\alpha_{j}<i_{1}<i_{2}<j_{2}=\alpha_{j}+3$. It follows that $W_{n}[F_{i}^{c}]$ is connected, a contradiction. So $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Moreover, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m-1$, we get $\alpha_{i}=m$. It follows that $W_{n}[F_{i}^{c}]$ is connected, again a contradiction. Hence our assumption that $j_{1}\leq i_{1}<i_{2}\leq j_{2}$ is false.
@@
    Suppose $i_{1}<j_{1}$. Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. First, we show that $F^{\prime}\in M(\Delta_{3}(W_{n}))$. We have $i_{1}<j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}=\alpha_{j}-2\geq 2$ (as $\alpha_{j}\geq 4$). Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-1$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
    Now, suppose that $i_{1}\geq j_{1}$. If $i_{2}\leq j_{2}$, then $j_{1}\leq i_{1}<i_{2}\leq j_{2}$, which is a contradiction. So $i_{2}>j_{2}$. Let $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Since $\alpha_{j}\geq 4$, we have $j_{1}\geq 3$. Then $3\leq j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}<i_{2}$. It follows that $i_{2}\nsim j_{1}$ and $i_{2}\nsim\alpha_{j}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $j_{1}=\alpha_{j}-1$ and $i_{2}\geq\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (iii) and $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).

### v2

@@ 3.1. Shelling Order
    If $\alpha_{j}\geq n-4$, then $j_{1},j_{2}>\alpha_{j}$ implies $C_{n}^{2}[F_{j}^{c}]$ is connected. So $\alpha_{j}\leq n-5$, and hence $\alpha_{i}\geq 2m-\alpha_{j}\geq 5$. We have $j_{2}>j_{1}>\alpha_{j}>m\geq\alpha_{i}\geq 5$. Thus $j_{2}\nsim\alpha_{i}$. Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. Now, if $j_{1}\nsim\alpha_{i}$, then $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{1}\nsim\alpha_{i}$ and $j_{2}\nsim\alpha_{j}$ imply that $j_{2}\nsim\alpha_{i}$).
    We now assume that $j_{1}\sim\alpha_{i}$. Suppose $j_{2}\sim\alpha_{i}$. Then $j_{1}<\alpha_{i}\leq m<\alpha_{j}<j_{2}$ implies that either $\alpha_{i}=m$, $\alpha_{j}=m+1$, $j_{2}=m+2$ and $j_{1}\in\{m-2,m-1\}$, or $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. In the former case, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}=m-2=\alpha_{j}-3$. This means that $F_{j}$ satisfies $(\mathcal{D}_{1})$, a contradiction. So, we have $\alpha_{i}=1$, $j_{2}=n-1$, and $j_{1}=0$. Then $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ and $\alpha_{j}>m$ implies that $\alpha_{j}=n-1=j_{2}$, which is again a contradiction. Therefore, $j_{2}\nsim\alpha_{i}$. Now, if $j_{1}\nsim j_{2}$, then $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$.
    So, let $j_{1}\sim j_{2}$. Then $F_{q}\notin M(\Delta_{3}(C_{n}^{2}))$. Let $F^{\prime c}:=\{\alpha_{i}\}\sqcup\{\alpha_{j},j_{2}\}$. Since $C_{n}^{2}[F_{j}^{c}]$ is disconnected and $j_{1}\sim j_{2}$, we get $j_{2}\nsim\alpha_{j}$. Therefore, $j_{2}\nsim\alpha_{i}$ implies that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Further, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$, we get $F^{\prime}\prec F_{j}$ by Proposition 3.6. Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
We show that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$, and thus $F_{r}=F_{q}$ satisfies ($\ast$3.1).
@@
    Since $m<\alpha_{i}<\alpha_{j}$, we have $\alpha_{j}\geq m+2$. Then $\alpha_{j}<_{\mathcal{O}}j_{2}<\alpha_{j}$ implies that $j_{2}\leq m-3$. It follows that $j_{1}<j_{2}\leq m-3<m<\alpha_{i}<\alpha_{j}$. Thus $j_{2}\nsim\alpha_{i}$. Further, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    We have $j_{2}>j_{1}>\alpha_{j}>\alpha_{i}>m>1$. So $j_{2}\nsim\alpha_{i}$. Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, $j_{1}\nsim j_{2}$ or $j_{1}\nsim\alpha_{j}$. Hence $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{1}\nsim\alpha_{j}$ implies that $j_{1}\nsim\alpha_{i}$).
@@
    Since $m<\alpha_{i}<\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}j_{1}<\alpha_{j}$, we get $j_{1}\leq m-3$. Hence $j_{1}\leq m-3<m<\alpha_{i}<\alpha_{j}<j_{2}$. Thus $j_{1}\nsim\alpha_{i}$. Now, since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, we have $j_{1}\nsim j_{2}$ or $j_{2}\nsim\alpha_{j}$. This implies that $F_{q}\in M(\Delta_{3}(C_{n}^{2}))$ (as $j_{2}\nsim\alpha_{j}$ implies that $j_{2}\nsim\alpha_{i}$).
@@
Suppose that $F_{j}\in\mathcal{D}$. Then there exists an $F_{r}\in M(\Delta_{3}(C_{n}^{2}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
We have ${F_{i}}^{c}=\{\alpha_{i}\}\sqcup\{i_{1},i_{2}\}$, ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$ and $F_{i}\prec F_{j}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, i.e., $i_{0}=j_{0}$, then we have $\alpha_{i}=\alpha_{j}$. In this case, if $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then $F_{r}=F_{i}$ satisfies ($\ast$3.1). So we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$ whenever $F_{i}\in\mathcal{T}_{j_{0}}$. Moreover, if $F_{i}\notin\mathcal{T}_{j_{0}}$, then $F_{i}\prec F_{j}$ and $F_{j}\in\mathcal{D}$ imply that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}$ by by Definition 3.4 (ii) and Definition 3.5. Note that since $n\geq 9$, we have $45\leq m\leq n-$ 4. Now, $F_{j}\in\mathcal{D}$ implies that $F_{j}$ satisfies one of the conditions from $(\mathcal{D}_{1})$ to $(\mathcal{D}_{4})$.
@@
    Let Let $F^{\prime c}:=(F_{j}^{c}\setminus\{j_{2}\})\sqcup\{i_{1}\}$ and $F^{\prime\prime c}:=(F_{j}^{c}\setminus\{j_{2}\})\sqcup\{i_{2}\}$. First, we show that if $i_{1}<j_{1}$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) and if $i_{2}>j_{2}$, then $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). Later, we discuss the case when $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$.
    First, suppose $i_{1}<j_{1}$. Since $\alpha_{j}<_{\mathcal{O}}j_{1}$, $j_{1}<\alpha_{j}$ and $\alpha_{j}>m$, we get $j_{1}<2m-\alpha_{j}$ by Remark 3.1 (ii), and hence $i_{1}<2m-\alpha_{j}$, thereby implying that $\alpha_{j}<_{\mathcal{O}}i_{1}$. This means that $F^{\prime c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$. Since $i_{1}<j_{1}=\alpha_{j}-3<\alpha_{j}\leq n-3$, we get $\alpha_{j}\nsim i_{1}$ and $\alpha_{j}\nsim j_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$. Therefore, $F^{\prime}\prec F_{j}$ by Definition 3.5 (iii) and (iv). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
    Let $i_{2}>j_{2}$. We have $i_{2}>j_{2}>\alpha_{j}>m$. Therefore $\alpha_{j}<_{\mathcal{O}}i_{2}$, and thus, $\alpha_{j}<_{\mathcal{O}}j_{1}$ and $j_{1}<\alpha_{j}<i_{2}$ implies that $F^{\prime\prime c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Since $\alpha_{j}\geq 6$, $j_{1}=\alpha_{j}-3\geq 3$. So we have $3\leq j_{1}=\alpha_{j}-3<\alpha_{j}<\alpha_{j}+1=j_{2}<i_{2}$. It follows that $j_{1}\nsim\alpha_{j}$ and $j_{1}\nsim i_{2}$. Hence $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $j_{1}=\alpha_{j}-3$ and $i_{2}\geq\alpha_{j}+2$, $F^{\prime\prime}\notin\mathcal{D}$. Therefore $F^{\prime\prime}\prec F_{j}$ by Definition 3.5 (iii), and thus $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
    Now, suppose $i_{1}\geq j_{1}$ and $i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then $\alpha_{j}=\alpha_{i}\neq i_{2}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. Thus, $i_{2}\leq j_{2}=\alpha_{j}+1$ implies that $i_{2}<\alpha_{j}$. Therefore, we have $\alpha_{j}-3=j_{1}<i_{1}<i_{2}<\alpha_{j}$. It follows that $C_{n}^{2}[F_{i}^{c}]$ is connected, a contradiction. Hence $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We know that $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m+1$. Thus, $\alpha_{i}\in\{m-1,m\}$. We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Therefore, since $C_{n}^{2}[F_{i}^{c}]$ is disconnected, we get $\alpha_{i}=m-1$, $i_{1}=j_{1}=m-2$ and $i_{2}=j_{2}=m+2$. Hence $F_{r}=F_{i}$ satisfies ($\ast$3.1).
@@
    Suppose $j_{1}\leq i_{1}<i_{2}\leq j_{2}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then we have $\alpha_{j}=\alpha_{i}\neq i_{1}$ and $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$. This implies that $\alpha_{j}<i_{1}<i_{2}<j_{2}=\alpha_{j}+3$. It follows that $C_{n}^{2}[F_{i}^{c}]$ is connected, a contradiction. So $F_{i}\notin\mathcal{T}_{j_{0}}$.
    We have $m-2=j_{1}\leq i_{1}<i_{2}\leq j_{2}=m+2$. Moreover, since $\alpha_{i}<_{\mathcal{O}}\alpha_{j}=m-1$, we get $\alpha_{i}=m$. It follows that $C_{n}^{2}[F_{i}^{c}]$ is connected, again a contradiction. Hence our assumption that $j_{1}\leq i_{1}<i_{2}\leq j_{2}$ is false.
@@
    Suppose $i_{1}<j_{1}$. Let $F^{\prime c}:=(F_{j}^{c}\setminus\{j_{1}\})\sqcup\{i_{1}\}$. We have $i_{1}<j_{1}<\alpha_{j}<m$. Therefore $\alpha_{j}<_{\mathcal{O}}i_{1}$, and thus, $\alpha_{j}<_{\mathcal{O}}j_{2}$ and $i_{1}<\alpha_{j}<j_{2}$ implies that $F^{\prime c}=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. First, we show that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. We have $i_{1}<j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}$. So $j_{2}\nsim\alpha_{j}$. Now, if $i_{1}\nsim\alpha_{j}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. So, let $i_{1}\sim\alpha_{j}$. Then $i_{1}=\alpha_{j}-2\geq 2$ (as $\alpha_{j}\geq 4$). Thus $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Further, since $i_{1}<j_{1}=\alpha_{j}-1$, $F^{\prime}\notin\mathcal{D}$. Thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (iii). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
    Now, suppose that $i_{1}\geq j_{1}$. If $i_{2}\leq j_{2}$, then $j_{1}\leq i_{1}<i_{2}\leq j_{2}$, which is a contradiction. So $i_{2}>j_{2}$. Let $F^{\prime\prime c}:=(F_{j}^{c}\setminus\{j_{2}\})\sqcup\{i_{2}\}$. Since $\alpha_{j}<_{\mathcal{O}}j_{2}$, $j_{2}>\alpha_{j}$ and $\alpha_{j}<m$, we get $j_{2}\geq 2m-\alpha_{j}$ by Remark 3.1 (i), and hence $i_{2}>2m-\alpha_{j}$, thereby implying that $\alpha_{j}<_{\mathcal{O}}i_{2}$. This means that $F^{\prime\prime c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$. Since $\alpha_{j}\geq 4$, we have $j_{1}\geq 3$. Then $3\leq j_{1}=\alpha_{j}-1<\alpha_{j}<\alpha_{j}+3=j_{2}<i_{2}$. It follows that $i_{2}\nsim j_{1}$ and $i_{2}\nsim\alpha_{j}$. Hence, $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $j_{1}=\alpha_{j}-1$ and $i_{2}\geq\alpha_{j}+4$, $F^{\prime\prime}\notin\mathcal{D}$. Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.5 (iii) and $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1).